# Catalog of Mock Stellar Streams in Milky Way-like Galaxies

### v1

@@ Abstract
Dynamically cold stellar streams from tidally dissolved globular clusters (GCs) serve as excellent tools to measure the Galactic mass distribution and show promise to probe the nature of dark matter. For successful application of these tools to observations, it is essential to have an accurate model of stellar stream properties on the Galactic scale. To this end we produce a mock catalog of stellar streams in four simulated Milky Way-like galaxies. We build the catalog with three main components: a model for the formation and disruption of globular clusters based on cosmological simulations, time-dependent potentials constructed with basis function expansions for integrating stream orbits, and an improved particle spray algorithm for efficient generation of stellar streams. We find that the observable widths and lengths of mock streams as a function of galactocentric radius are well described by power-laws for streams beyond 10 kpc. We generate mock photometry for Gaia, LSST, and Roman, and find that the latter two surveys will increase the number of observable stars in GC stellar streams by several orders of magnitude. Our full catalog, containing stream populations across four different galaxy realizations, is publicly available and can be used to study stream population statistics and to calibrate models which use stellar streams to understand our Galaxy.
@@ 1 Introduction
Globular clusters (GCs) are some of the best available tools for understanding the assembly history and evolution of the Milky Way galaxy (MW). Since most clusters are observed to have an age $\sim 10$ Gyr (Krauss & Chaboyer, 2003; Dotter et al., 2011), they have been orbiting our Galaxy for at least a few Gyr if accreted onto the MW (ex-situ formed clusters) or longer if born in the central part of the MW (in-situ formed clusters). As GCs orbit the Galaxy, they are subject to tidal forces that cause some stars to leave the cluster and form an unbound stellar stream. Recently, the Gaia mission (Gaia Collaboration et al., 2018, 2023) has significantly expanded our knowledge of MW streams (Bonaca & Price-Whelan, 2025), of which $\sim$80 are thought to originate from GCs (Mateu, 2023). Since the velocity dispersion of unbound GC stars is smaller than the typical orbital velocity around the MW, GC streams are dynamically cold, retaining signatures of their tidal histories for up to several Gyr.
@@
One of the most exciting applications which shows promise to be better understood through stellar streams is constraining the mass distribution within the Galaxy (Johnston et al., 1999; Helmi et al., 1999; Koposov et al., 2010; Sanders & Binney, 2013; Bonaca & Hogg, 2018; Reino et al., 2021; Ibata et al., 2024; Nibauer & Bonaca, 2025). This is typically done by assuming an analytic form of the gravitational potential and then varying the potential’s free parameters and/or symmetry constraints to identify which model can best recreate a given stream or select group of streams. In addition, Nibauer & Bonaca (2025) showed that the GD-1 stream (Grillmair & Dionatos, 2006) can be used to constrain the potential by direct differentiation of the stream track in phase space, without first having to assume an analytic form for the potential. Stream modeling can also be used to infer properties of satellite galaxies interacting with the MW. For example, Shipp et al. (2021) used measurements from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016), the Southern Stellar Stream Spectroscopic Survey ($S^{5}$; Li et al., 2019), and Gaia to estimate the mass of the Large Magellanic Cloud (LMC) by modeling its perturbations on different stellar streams, obtaining a value of $1.88^{+0.35}_{-0.40}\times 10^{11}\,{\mathrm{M}_{\odot}}$. Vasiliev et al. (2021) performed a similar type of analysis using restricted N-body modeling to recreate Gaia measurements of the Sagittarius stream (Ibata et al., 2001) in the presence of the LMC, computing a MW virial mass of $(9.0\pm{1.3})\times 10^{11}\,{\mathrm{M}_{\odot}}$.
@@
The density of stars within streams may exhibit variations caused by gravitational perturbations from substructure in the Galactic potential, such as a tilted stellar disk (Nibauer et al., 2024), giant molecular clouds (Amorisco et al., 2016), the Galactic bar (Pearson et al., 2017; Bonaca et al., 2020), and dark matter subhaloes (Johnston et al., 2002; Ibata et al., 2002; Carlberg, 2012; Erkal et al., 2016; Banik et al., 2018). Disentangling the baryonic perturbations from those potentially caused by dark matter (DM) is key in the hope of using stellar streams to constrain the lower end of the DM subhalo mass function. Detecting streams at large galactocentric radii will be extremely important for this purpose. A key effort in this direction is to model streams beyond the stellar disk ($r_{\rm gal}\sim 15\rm\penalty 10000\ kpc$), both in terms of physical properties and prospects for observing them.
@@
Many published works model streams via the particle-spray method, which releases stream particles near the Lagrange points at frequent timesteps with small offsets in their positions and velocities (Varghese et al., 2011; Lane et al., 2012; Küpper et al., 2012; Bonaca et al., 2014; Gibbons et al., 2014; Fardal et al., 2015; Roberts et al., 2025). Integrating the orbits of these particles to the present day produces an elongated stream profile, generally consistent with the observations, providing a reasonably accurate method which is significantly faster than full N-body modeling methods. Recently, Chen et al. (2025b) developed an improved particle-spray algorithm where the phase-space offsets of released particles follow a multivariate Gaussian distribution. The new method matches the morphology and action space distribution of streams from N-body simulations with an error of only 10%. This method allows computationally efficient modeling of galactic stream populations with greater detail than previously possible.
@@
In this paper, we build upon the work of Pearson et al. (2024) and generate mock GC stream populations in four simulated MW-like galaxies with realistic time-evolving potentials. We use the Chen et al. (2025b) particle spray algorithm to integrate the orbits of unbound stars for several Gyr. We analyze overall properties of the final stream populations and study how facilities such as LSST and the future Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2013, 2015) will expand our ability to detect Galactic GC streams. We also make our stream catalog publicly available to be used as a tool for analyzing expected trends in the stream populations, and for calibration of models which attempt to understand the Galaxy through the use of GC streams.
@@ 2 Model Setup
In this section, we describe the details of the modeling used to generate our mock stream catalog. The key components are the initialization of the GC population, dynamical evolution of the GCs, and the generation of streams via the particle spray method.
@@ 2.1 Globular Cluster Evolution
The modeling of GC evolution requires a detailed understanding of not only the birth of clusters, both in-situ and ex-situ, but also their mass loss over time through stellar evolution and tidal disruption. For these purposes, we employ the GC evolution model by Chen & Gnedin (2022), which is further developed in Chen & Gnedin (2024a). The model post-processes snapshots of a selected cosmological simulation and creates catalogs of the GC population, keeping track of cluster formation and evolution. The model produces key observables including mass, metallically, positions, and velocities of individual GCs. The model has been applied to a wide range of galaxies in cosmological simulations and can successfully match the observed properties of the Galactic GC system.
A full description of the most recent version of the model is given in Chen & Gnedin (2024a). The three main features are: cluster formation, particle assignment for orbit tracking, and mass loss through tidal disruption along the GC’s orbit. GC formation is triggered when the host galaxy mass accretion rate exceeds a specified threshold, which is an adjustable parameter of the model. The total mass of the cluster population formed at a given snapshot epoch is proportional to the cold gas mass of the parent galaxy, following Kravtsov & Gnedin (2005). Given this population, the individual masses of GCs are sampled from a Schechter (1976) initial cluster mass function with a lower bound of $10^{4}\,{\mathrm{M}_{\odot}}$ and an exponential cutoff of $10^{7}\,{\mathrm{M}_{\odot}}$. To obtain kinematic information for the clusters at this epoch, the model assigns clusters to collisionless particles in the background cosmological simulation and uses their positions and velocities. The mass evolution takes into account both stellar evolution and tidal disruption along the cluster orbit. Once a cluster is initialized with mass $M_{0}$, the model treats the mass-loss due to stellar evolution as instantaneous:
@@
where the remaining fraction $\mu_{\rm sev}$ is taken to be $0.55$ (Gieles & Gnedin, 2023). This assumption is justified because the timescale for stellar evolution of high mass stars (tens of Myr) is much shorter than the typical age of GCs. Here $M_{\rm i}$ represents the initial mass of the cluster after stellar evolution but before any tidal disruption.
The stream is populated by stars unbound from the cluster by tidal forces. The most recent prescription for tidal mass-loss outlined in Chen & Gnedin (2024a) is given by:
@@
where $\Omega_{\mathrm{tid}}(t)$ is the effective tidal frequency, approximated via the first and third eigenvalues ($\lambda_{1},\lambda_{3}$) of the tidal tensor:
@@
The model calculates the tidal tensor using a second order finite difference method for snapshots where the potential for the host galaxy is provided. However, this only permits the masses of the GCs to be updated at every full snapshot. To improve the time resolution, we use a combination of the GC masses computed by the Chen & Gnedin (2024a) model as well as a higher-resolution calculation of the mass-loss rate outlined in Section 2.3.
We note that by default, the GC evolution model calculates the mass of each GC at every snapshot, but only outputs the mass to the user at the present day. To address this for the purposes of this work, we make a small modification of the code to have it output the cluster masses at every snapshot. Due to this modification, we re-run the model on the four cosmological halos utilized in this work. Both re-running the model as well as utilizing a different mass-loss calculation method results in our GCs being slightly different from the GCs in the published catalogs of Chen & Gnedin (2024a). These differences are explained in detail in Appendix C, but we stress that overall our GCs are still tailored to match the MW system in terms of the mass, metallicity, and radial distributions.
@@ 2.2 Background Cosmological Simulations
To generate a realistic model for an initial GC population in MW-like galaxies, we utilize two halos from the Illustris TNG50-1 (TNG) cosmological simulation (Nelson et al., 2019), and two halos from the Latte suite of FIRE-2 (FIRE) simulations (Wetzel et al., 2016). The TNG galaxies have subhalo IDs 523889 and 519311 in the TNG halo catalog, and the FIRE halos are m12i and m12w. Chen & Gnedin (2024b) identified these four halos to be a close match to the Milky Way for the purpose of applying their GC post-processing model. The criteria used for determining suitable galaxies in Chen & Gnedin (2024b) include having a present mass and circular velocity similar to that observed for the MW, no major merger (defined as having a mass ratio greater than 1:4) in the last 10 Gyr, one major merger between 10-12 Gyr ago representing the Gaia-Sausage/Enceladus merger (Belokurov et al., 2018; Helmi et al., 2018), and to have formed 25-35% of its stellar mass by 10 Gyr ago (Leitner, 2012). Our selected halos satisfy all these requirements and are suitable candidates to approximate the present day stream distribution of the MW.
@@
TNG and FIRE provide individual snapshots of halo particles, separated by $10^{8}-10^{9}$ yr, which by itself is not suitable for computing the orbits of millions of stream stars for several Gyr. For the actual computation of stream orbits, we construct analytical basis function expansion (BFE) potentials based on the TNG and FIRE snapshots using the agama software (Vasiliev, 2019). We construct the analytic time-dependent potentials in the following way: for each snapshot, we fit a BFE potential to the snapshot given the positions and masses of the stellar, gas, and dark matter particles. The dark matter particles are fit with a standard multipole expansion of spherical harmonics with $l_{\text{max}}$ taken to be $12$. The stellar and gas particles are fit with an azimuthal Fourier harmonic expansion with $m_{\text{max}}=6$. The method for determining these parameters and further details are outlined in Appendix A. We construct the time-variable potentials by linearly interpolating the expansion coefficients between the snapshots. For each simulation the overall time span is taken to be $z=0.3$ to the present day (about 3.5 Gyr). For the TNG halos, this corresponds to every snapshot starting from 78 to 99 (present day), which are spaced by intervals ranging from 120235120-235 Myr for TNG’s cosmological parameters. For FIRE, we use the same snapshots with full particle data as in Chen & Gnedin (2024b) from $z=0.3$ to present, corresponding to snapshots 446, 486, 534, and 590-600. The spacing between these snapshots respectively is 987 Myr, 1140 Myr, 1280 Myr, and 2.22 Myr for the remaining ten.
@@
For the FIRE halo m12i, we find that the principal axis of the disk tilts by $24^{\circ}$ over this time span, and therefore we modify the above prescription for this halo only and model its potential just with multipole expansion. This is because all orbits are computed in the galactocentric frame of the halo at the last snapshot, but the agama Fourier harmonic expansion fitting procedure requires the disk to be aligned with the XYX-Y plane to work properly. Since multipole expansion is not as effective at modeling flattened systems, this results in our disk potential for m12i not being as accurate as in the other three halos. However, we find that fitting one of our other galaxies with only multipoles results in about 2% average difference in the orbits compared to using both BFE functions. Since this difference is small, we still proceed to use the m12i halo, but users of the catalog should be aware of this if they choose to use the m12i streams.
@@
Instead of integrating orbits of particles immediately from cluster birth (or from time of accretion onto the MW), we select a lookback time to begin tracking the orbits of tidally stripped stars from each cluster. This is because phase mixing causes the orbits of evaporated stars to deviate significantly from the progenitor with time, limiting the duration of coherent streams on most orbits to a few Gyr. Moreover, Chen et al. (2025a) showed that the duration of most observed GC stellar streams are $\leq 1\rm\penalty 10000\ Gyr$ (although some streams are likely to be much older – see Erkal et al. 2016). It is thus unnecessary to include these early-released particles in our catalog. Based on these arguments we take our initial snapshot for orbit tracking of evaporated particles to be $z=0.3$.

### v2

@@ Abstract
Dynamically cold stellar streams from tidally dissolved globular clusters (GCs) serve as excellent tools to measure the Galactic mass distribution and show promise to probe the nature of dark matter. For successful application of these tools to observations, it is essential to have models of stellar stream properties on the Galactic scale. To this end we produce a mock catalog of stellar streams in four Milky Way-mass galaxies from cosmological simulations. We build the catalog with three main components: a model for the formation and disruption of globular clusters based on cosmological simulations, time-dependent potentials constructed with basis function expansions for integrating stream orbits, and an improved particle-spray algorithm for efficient generation of stellar streams. We generate mock photometry for Gaia, LSST, and Roman, and find that the latter two surveys will increase the number of observable stars in GC stellar streams by several orders of magnitude, with LSST alone being able to observe 1000 or more stars in 40-100 streams beyond the stellar disk. In addition, we find that the observable widths and lengths of mock streams as a function of galactocentric radius are well described by power-laws with slopes of 1.8-1.8 and 2-2 for streams beyond 10 kpc. Our full catalog, containing stream populations across four different galaxy realizations, is publicly available and can be used to study stream population statistics and to calibrate models which use stellar streams to understand our Galaxy.
@@ 1 Introduction
Globular clusters (GCs) are some of the best available tools for understanding the assembly history and evolution of the Milky Way galaxy (MW). Since most clusters are observed to have an age $\sim 10$ Gyr (Krauss and Chaboyer, 2003; Dotter et al., 2011), they have been orbiting our Galaxy for at least a few Gyr if accreted onto the MW (ex-situ formed clusters) or longer if born in the central part of the MW (in-situ formed clusters). While GC stars can become unbound through internal dynamics such as two-body relaxation (Spitzer, 1987; Meylan and Heggie, 1997) and ejection from black hole interactions (Gieles and Gnedin, 2023; Roberts et al., 2025), the dominate mechanism is through tidal heating (Gnedin and Ostriker, 1997; Gnedin et al., 1999; Baumgardt and Makino, 2003), which can cause enough mass-loss to form an unbound stellar stream. Recently, the Gaia mission (Gaia Collaboration et al., 2018, 2023) has significantly expanded our knowledge of MW streams (Bonaca and Price-Whelan, 2025), of which $\sim$80 are thought to originate from GCs (Mateu, 2023). Since the velocity dispersion of unbound GC stars is smaller than the typical orbital velocity around the MW, GC streams are dynamically cold, retaining signatures of their tidal histories for up to several Gyr. This makes them ideal for studying a wide range of gravitational physics within the MW itself, both on small and large scales.
@@
One of the most exciting applications which shows promise to be better understood through stellar streams is constraining the mass distribution within the Galaxy (Johnston et al., 1999; Helmi et al., 1999; Koposov et al., 2010; Sanders and Binney, 2013; Bonaca and Hogg, 2018; Reino et al., 2021; Ibata et al., 2024; Nibauer and Bonaca, 2025). This is typically done by assuming an analytic form of the gravitational potential and then varying the potential’s free parameters and/or symmetry constraints to identify which model can best recreate a given stream or select group of streams. In addition, Nibauer and Bonaca (2025) showed that the GD-1 stream (Grillmair and Dionatos, 2006) can be used to constrain the potential by direct differentiation of the stream track in phase space, without first having to assume an analytic form for the potential. Stream modeling can also be used to infer properties of satellite galaxies interacting with the MW. For example, Shipp et al. (2021) used measurements from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016), the Southern Stellar Stream Spectroscopic Survey ($S^{5}$; Li et al., 2019), and Gaia to estimate the mass of the Large Magellanic Cloud (LMC) by modeling its perturbations on different stellar streams, obtaining a value of $1.88^{+0.35}_{-0.40}\times 10^{11}\,{\mathrm{M}_{\odot}}$. Vasiliev et al. (2021) performed a similar type of analysis to recreate Gaia measurements of the Sagittarius stream (Ibata et al., 2001) in the presence of the LMC, computing a MW virial mass of $(9.0\pm{1.3})\times 10^{11}\,{\mathrm{M}_{\odot}}$.
@@
In addition to modeling the global potential of the MW and its satellites, streams are also subject to perturbations from dark matter subhalos (e.g., Johnston et al., 2002; Ibata et al., 2002; Carlberg, 2012; Erkal et al., 2016; Banik et al., 2018) which can in principle be used to constrain the subhalo mass function. However, the density of stars within streams may also exhibit variations caused by baryonic perturbations from substructure in the Galactic potential, such as a tilted stellar disk (Nibauer et al., 2024), giant molecular clouds (Amorisco et al., 2016), and the Galactic bar (Pearson et al., 2017; Bonaca et al., 2020). Epicyclic overdensities can also appear as density variations in the stream track without a perturbation Küpper et al. (2008, 2010). Disentangling these additional perturbations from those potentially caused by dark matter is key in the hope of using stellar streams to constrain the dark matter subhalo mass function. Detecting streams at large galactocentric radii will be extremely important for this purpose, where interactions with baryonic components of the potential are less likely. A key effort in this direction is to model streams beyond the stellar disk ($r_{\rm gal}\sim 15\rm\penalty 10000\ kpc$), both in terms of physical properties and prospects for observing them.
In order to model these GC streams far out from the Galactic center, one must first have an understanding of how the GC population itself is evolving over time for a given model Galaxy, as well as a method for generating streams from the clusters based on their mass-loss rates along a given orbit. This was done in Panithanpaisal et al. (2025), which applied the GC formation and evolution models from Grudić et al. (2023) and Rodriguez et al. (2023) to study two GC streams evolved in a MW-like Galaxy from the FIRE-2 cosmological simulation suite (Wetzel et al., 2016; Hopkins et al., 2018; Wetzel et al., 2023). Looking at the real MW GC population, Ferrone et al. (2023) modeled the theoretical extra-tidal features of MW GCs for which 6D phase space coordinates are known and found coherent stream-like structures should be present as far out as $300$ kpc from the Galactic center.
@@
In studies such as these which model many stream realizations in a realistic Galactic potential, an approximation to direct N-body is needed to speed up the computation time. For this purpose, many works utilize the particle-spray method, which releases stream particles near the Lagrange points at frequent timesteps with small offsets in their positions and velocities (Varghese et al., 2011; Lane et al., 2012; Küpper et al., 2012; Bonaca et al., 2014; Gibbons et al., 2014; Fardal et al., 2015; Roberts et al., 2025). Integrating the orbits of these particles to the present day produces an elongated stream profile, generally consistent with the observations, providing a reasonably accurate method which is significantly faster than full N-body modeling methods. Recently, Chen et al. (2025b) developed an improved particle-spray algorithm where the phase-space offsets of released particles follow a multivariate Gaussian distribution. The new method matches the morphology and action space distribution of streams from N-body simulations with an error of only 10%. This method allows computationally efficient modeling of Galactic stream populations with greater detail than previously possible.
@@
In this paper, we build upon the work of Pearson et al. (2024) and generate mock GC stream populations in four simulated MW-like galaxies with realistic time-evolving potentials. We use the Chen et al. (2025b) particle-spray algorithm to integrate the orbits of unbound stars for several Gyr. We analyze overall properties of the final stream populations and study how facilities such as LSST and the future Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2013, 2015) will expand our ability to detect Galactic GC streams. We also make our stream catalog publicly available to be used as a tool for analyzing expected trends in the stream populations, and for calibration of models which attempt to understand the Galaxy through the use of GC streams.
@@ 2 Model Setup
In this section, we describe the details of the modeling used to generate our mock stream catalog. The key components are the initialization of the GC population, dynamical evolution of the GCs, and the generation of streams via the particle-spray method.
@@ 2.1 Globular Cluster Evolution
The modeling of GC evolution requires a detailed understanding of not only the birth of clusters, both in-situ and ex-situ, but also their mass loss over time through stellar evolution and tidal disruption. For these purposes, we employ the GC evolution model by Chen and Gnedin (2022), which is further developed in Chen and Gnedin (2024a). The model post-processes snapshots of a selected cosmological simulation and creates catalogs of the GC population, keeping track of cluster formation and evolution. The model produces key observables including mass, metallically, positions, and velocities of individual GCs. The model has been applied to a wide range of galaxies in cosmological simulations and can successfully match the observed properties of the Galactic GC system.
A full description of the most recent version of the model is given in Chen and Gnedin (2024a). The three main features are: cluster formation, particle assignment for orbit tracking, and mass loss through tidal disruption along the GC’s orbit. GC formation is triggered when the host galaxy mass accretion rate exceeds a specified threshold, which is an adjustable parameter of the model. The total mass of the cluster population formed at a given snapshot epoch is proportional to the cold gas mass of the parent galaxy, following Kravtsov and Gnedin (2005). Given this population, the individual masses of GCs are sampled from a Schechter (1976) initial cluster mass function with a lower bound of $10^{4}\,{\mathrm{M}_{\odot}}$ and an exponential cutoff of $10^{7}\,{\mathrm{M}_{\odot}}$. To obtain kinematic information for the clusters at this epoch, the model assigns clusters to collisionless particles in the background cosmological simulation and uses their positions and velocities. The mass evolution takes into account both stellar evolution and tidal disruption along the cluster orbit. Once a cluster is initialized with mass $M_{0}$, the model treats the mass-loss due to stellar evolution as instantaneous:
@@
where the remaining fraction $\mu_{\rm sev}$ is taken to be $0.55$ (Gieles and Gnedin, 2023). This assumption is justified because the timescale for stellar evolution of high mass stars (tens of Myr) is much shorter than the typical age of GCs. Here $M_{\rm i}$ represents the initial mass of the cluster after stellar evolution but before any tidal disruption.
The stream is populated by stars unbound from the cluster by tidal forces. The most recent prescription for tidal mass-loss outlined in Chen and Gnedin (2024a) is given by:
@@
where $\Omega_{\mathrm{tid}}(t)$ is the effective tidal frequency, approximated via the first and third eigenvalues ($\lambda_{1},\lambda_{3}$) of the tidal tensor of the host potential:
@@
The model calculates the tidal tensor using a second order finite difference method for snapshots where the potential for the host galaxy is provided. However, this only permits the masses of the GCs to be updated at every full snapshot. To improve the time resolution, we use a combination of the GC masses computed by the Chen and Gnedin (2024a) model as well as a higher-resolution calculation of the mass-loss rate outlined in Section 2.3. Our mass-loss rate model is calibrated from N-body models ran in Gieles and Gnedin (2023) and the parameters were optimized in Chen and Gnedin (2024a) to be able to match the MW in terms of total GC number, present day mass function, radial distribution, metallicity distribution, and velocity dispersion. While other phenomena can affect individual GC mass-loss rates that are either not incorporated or may be unresolved in Equation 2 such as GC concentration (Baumgardt and Makino, 2003), heating from black holes (Gieles and Gnedin, 2023), and tidal shocks (e.g., Gnedin et al., 1999), statistically our model matches the present day trends of the total MW GC population.
We note that by default, the GC evolution model calculates the mass of each GC at every snapshot, but only outputs the mass to the user at the present day. To address this for the purposes of this work, we make a small modification of the code to have it output the cluster masses at every snapshot. Due to this modification, we re-run the model on the four cosmological halos utilized in this work. Both re-running the model as well as utilizing a different mass-loss calculation method results in our GCs being slightly different from the GCs in the published catalogs of Chen and Gnedin (2024a). These differences are explained in detail in Appendix A, but we stress that overall our GCs are still tailored to match the MW system in terms of the mass, metallicity, and radial distributions.
@@ 2.2 Background Cosmological Simulations
To generate a realistic model for an initial GC population in MW-like galaxies, we utilize two halos from the Illustris TNG50-1 (TNG) cosmological simulation (Nelson et al., 2019; Pillepich et al., 2019), and two halos from the Latte suite of FIRE-2 (FIRE) simulations (Wetzel et al., 2016; Hopkins et al., 2018; Wetzel et al., 2023). The TNG galaxies have subhalo IDs 523889 and 519311 in the TNG halo catalog, and the FIRE halos are m12i and m12w. Chen and Gnedin (2024b) identified these four halos to be a close match to the Milky Way for the purpose of applying their GC post-processing model. The criteria used for determining suitable galaxies in Chen and Gnedin (2024b) include having a present mass and circular velocity similar to that observed for the MW, no major merger (defined as having a mass ratio greater than 1:4) in the last 10 Gyr, one major merger between 10-12 Gyr ago representing the Gaia-Sausage/Enceladus merger (Belokurov et al., 2018; Helmi et al., 2018), and to have formed 25-35% of its stellar mass by 10 Gyr ago (Leitner, 2012). Our selected halos satisfy all these requirements and are suitable candidates to approximate the present day stream distribution of the MW.
@@
TNG and FIRE provide individual snapshots of halo particles, separated by 0.110.1-1 Gyr, which by itself is not suitable for computing the orbits of millions of stream stars for several Gyr. For the actual computation of stream orbits, we construct analytical basis function expansion (BFE) potentials based on the TNG and FIRE snapshots using the agama software (Vasiliev, 2019). We construct the analytic time-dependent potentials in the following way: for each snapshot, we fit a BFE potential to the snapshot given the positions and masses of the stellar, gas, and dark matter particles. The dark matter particles are fit with a standard multipole expansion of spherical harmonics with $l_{\text{max}}$ taken to be $12$. The stellar and gas particles are fit with an azimuthal Fourier harmonic expansion with $m_{\text{max}}=6$. The method for determining these parameters and further details are outlined in Appendix B. We construct the time-variable potentials by linearly interpolating the expansion coefficients between the snapshots. For each simulation the overall time span is taken to be $z=0.3$ to the present day (about 3.5 Gyr). For the TNG halos, this corresponds to every snapshot starting from 78 to 99 (present day), which are spaced by intervals ranging from 120235120-235 Myr for TNG’s cosmological parameters. For FIRE, we use the first data release snapshots with full particle data from $z=0.3$ to present, corresponding to snapshots 446, 486, 534, and 590-600. The spacing between these snapshots respectively is 987 Myr, 1140 Myr, 1280 Myr, and 2.22 Myr for the remaining ten.
@@
For the FIRE halo m12i, we find that the principal axis of the disk tilts by $24^{\circ}$ over this time span, and therefore we modify the above prescription for this halo only and model its potential just with multipole expansion. This is because all orbits are computed in the galactocentric frame of the halo at the last snapshot, but the agama Fourier harmonic expansion fitting procedure requires the disk to be aligned with the XYX-Y plane to work properly. Since multipole expansion is not as effective at modeling flattened systems, this results in our disk potential for m12i not being as accurate as in the other three halos. However, we find that after fitting one of our other galaxies with only multipoles, the median separation between our reconstructed orbit to the true orbit at the last snapshot increases by only about 2% compared to using both BFE functions. Since this difference is small, we still proceed to use the m12i halo, but users of the catalog should be aware of this if they choose to use the m12i streams.
Our method of potential reconstruction through interpolation of BFE potentials has been shown to be an effective way for approximating the potentials of galaxies from cosmological simulations (e.g., Lowing et al., 2011; Sanders et al., 2020; Arora et al., 2024). In Appendix C, we show some example orbits of our analytic BFE potentials compared to the orbits of the GC particles in their respective cosmological simulations. In general, this prescription of potential modeling allows us to reliably recover the pericenter and apocenter values of the GC particles over our integration time. As other works have already shown, we find the method is less effective on orbits whose apocenters are extremely close to the Galactic center, $\lesssim 5$ kpc. However, these orbits are less likely to have kinematically cold streams to begin with. We also find that for some GCs with shorter periods, the reconstructed orbit becomes out of phase with the true one as small errors accumulate over many orbits. (Arora et al., 2024) found specifically that pericenters and apocenters can be recovered within $10\%$ over a baseline of 4 Gyr. While they use a higher temporal cadence than what is utilized in this work, we show in Appendix C that our method also reasonably recovers the period, pericenter, and apocenter of most orbits over a similar integration time, which are key quantities in determining the overall stream morphology.
@@
Instead of integrating orbits of stream particles immediately from cluster birth (or from time of accretion onto the MW), we select a lookback time to begin tracking the orbits of tidally stripped stars from each cluster. This is because phase mixing causes the orbits of evaporated stars to deviate significantly from the progenitor with time, limiting the duration of coherent streams on most orbits to a few Gyr. Moreover, Chen et al. (2025a) showed that the duration of most observed GC stellar streams are $\leq 1\rm\penalty 10000\ Gyr$ (although some streams are likely to be much older – see Erkal et al. 2016). It is thus unnecessary to include these early-released particles in our catalog. Based on these arguments we take our initial snapshot for orbit tracking of evaporated particles to be $z=0.3$.